### v1

Abstract
Magnetic reconnection in magnetized wakes of cosmic strings results in the release of a large amount of energy. This energy is released in a short period of time. In this work, we show that this sudden release of energy can result in a Gamma Ray Burst (GRB) of short duration. The magnetic reconnection occurs at several points of the cosmic string wake. The emerging shocks from these points have different velocities. These shocks will collide with each other and give rise to short bursts of energy. The emitted pulse of energy depends on the background magnetic field and the timescale associated with the magnetic reconnection in the cosmic string wake. We also obtain the synthetic lightcurve that occurs from multiple collisions of the shock waves that are emitted from the multiple points of magnetic reconnection in the wake region. Finally, we discuss about some current experimental evidence of short GRB’s which are in the same energy range as the ones predicted from our model.
cosmic string wakes, magnetic reconnection, GRBs.
I Introduction
Symmetry breaking phase transitions in the early universe give rise to topological defects. One of the topological defects that is predicted by these symmetry breaking phase transitions is the cosmic string defect. The cosmic string defect is a one dimensional defect generated due to a $U(1)$ symmetry breaking phase transition. Once formed, these strings will move through the early universe plasma. Depending on the exact nature of the phase transition, long strings or loops of cosmic strings are formed. Some of these defects are unstable and decay with the evolution of the universe, others are stable and evolve as a giant network of cosmic strings spanning the universe cosmicstrings . Theoreticians have predicted multiple observational signatures of these cosmic strings. Primary amongst these were the gravitational lensing of objects due to the conical space time of the cosmic string lensing . Other than the gravitational lensing, as the cosmic string moves through the plasma, it gives rise to wakes of overdensity behind it wakes . These wakes will lead to a distinct wedge shaped patterns in the 21 cm redshift survey maps branden1 ; dacunha . These wakes also give rise to temperature fluctuations which have been used for searches for these exotic objects by the Planck collaboration Planck_2016 . All these observational signatures are currently from density inhomogeneities generated as the cosmic string moves through the plasma.
Cosmic strings moving through the plasma have also been associated with the primordial magnetic fields in the early universe. Two cosmic strings moving past one another in the plasma lead to the generation of primordial magnetic field in the early universe vachaspati . The Harrison mechanism which leads to differential rotation between the heavier ions and the lighter electrons also leads to the generation of a primordial magnetic field close to the cosmic strings Harrison . In a recent work, we have shown that magnetic fields can be generated in cosmic string wakes due to the Biermann battery mechanism Sovan_2020 .
The wakes of these cosmic strings are long and narrow. The Biermann battery mechanism ohira produces magnetic fields that are oppositely directed along the two sides of the wake. A detailed study using the Sweet-Parker model has shown the possibility of magnetic reconnection in these magnetized string wakes Dilip_2023 . Magnetic reconnection occurs when two oppositely directed magnetic fields come very close to each other. It leads to the rearrangement of the magnetic field lines by breaking and cross-connecting to each other which results in a burst of energy. Numerical simulations using magnetohydrodynamic codes have also shown the evolution of the magnetic field and the formation of loops of magnetic field lines due to magnetic reconnection Soumen_2022 in cosmic string wakes.
Previously it has been predicted that magnetic reconnection can lead to a Gamma Ray Burst (GRB) Granot . As per the current literature, the different models of magnetic reconnection predict that the burst of energy due to magnetic reconnection is accompanied by the generation of shock waves yamada_review . In this work, we show that the collision of these shocks in the cosmic string can lead to a GRB. GRB’s are the bursts of energies observed by various telescopes. Their extreme luminosity and high energy output make them crucial for studying the early universe Paczynski_1998 . In this work, we obtain the possible GRB light pulse which will be generated from the magnetic reconnection in the cosmic string wake. As we will explain in detail, there is the possibility of multiple reconnection points occuring on the cosmic string wake and so we also obtain the synthetic GRB light curve based on our model of the light pulse.
The GRB’s are generally classified based on their luminosity and the time duration over which they occur piran . These are two of the basic features of any GRB. The short duration GRB’s last for less than two seconds while the long duration GRB’s usually last for more than two seconds. We find the GRB light curves which indicate that short duration GRB’s can be generated from magnetic reconnection in cosmic string wakes. We also find that the possible time durations and the fluence calculated in our model is within the range of the short duration GRB’s identified by satellites like BATSE (Burst And Transient Source Experiment) batse . Overall, we establish that magnetic reconnection in cosmic string wakes may provide a new way of identifying these exotic objects using both electromagnetic signals and signatures from density inhomogeneities.
II Magnetic Reconnection in String Wakes
In a recent work, we have shown that the small scale inhomogeneities which are perpendicular to the shock fronts in the cosmic string wakes generate magnetic fields in the wake due to the Biermann battery mechanism Sovan_2020 . Due to the symmetry of the wake, the fields generated are oppositely directed on the two sides of the wake. As the width of the wake is small, and the background particles stream into the wake overdensity, they push the magnetic lines together and magnetic reconnection occurs. A basic calculation based on the Sweet–Parker (SP) model Parker_1957 shows that a large amount of energy is released due to the magnetic reconnection in the cosmic string wake. The SP model approximated the problem as a two dimensional problem with a plasma of constant resistivity. The energy though large was released over a substantial period of time. Other models increased the reconnection rate by using a variable resistivity che . According to the SP model, the outflow velocity is dependent on the reconnection velocity. This velocity is given by, $V_{r}=\sqrt{{C_{H}\eta c}\over{4\pi l}}$. Here, $C_{H}=\frac{B_{0}}{\sqrt{4\pi\rho}}$ is the velocity with which the magnetic lines of force move towards each other, $\eta$ is the resistivity and $l$ is the length of the diffusion region, which is the region over which the magnetic reconnection takes place. The velocity $C_{H}$ is referred to as the hydromagnetic velocity. Shocks are generated at the points of reconnection and these shocks move in an outward direction away from the point of reconnection. The shocks are relativistic shocks and have different velocities and gamma factors. Since these shocks have different velocities, the shocks will collide with one another. In Fig.[1], we show the region in the cosmic string wake where the magnetic field lines come together and points of magnetic reconnection are formed.
Figure 1: Points of magnetic reconnection being formed in the cosmic string wake. Here X marks the position of the cosmic string.
At any given time $t_{i}$, the cosmic string wake has dimensions given by branden1 , $c_{1}t_{i}\times t_{i}v_{s}\gamma_{s}\times\delta\theta t_{i}v_{s}\gamma_{s}$ where, $c_{1}$ is of order $1$, $v_{s}$ is velocity of the moving string, $\gamma_{s}$ is the Lorentz factor associated with the velocity of the string. Since the string usually moves at relativistic velocities, the average velocity is given by $<v_{s}\gamma_{s}>$ $\sim 0.5$ dacunha . The $\delta\theta$ in the dimensions is the opening angle of the cosmic string wake. It is equal to the deficit angle of the space time around the cosmic string and is given by $\delta\theta\sim 4\pi G\mu$. The value of $G\mu$ is given by the symmetry breaking scale at which the cosmic string is formed vilenkin . Since this is very small (of the order of $10^{-7}$ or less), the width and the height of the wake is also very small compared to the length of the wake. The oppositely directed magnetic fields are along the $z-$direction so the relevant length would be along the $z-$direction. Usually a quasi 2-dimensional model is used to understand magnetic reconnection, the dimensions of the cosmic string wake also indicate that the two oppositely directed magnetic field are separated by a very small lengthscale.
As the string moves, the fluid motion due to the cosmic string metric pushes the fluid towards the center of the wake. The magnetic field lines also feel this hydrostatic pressure and are therefore pushed towards the center of the wake. As the magnetic field lines come very close to each other, points of magnetic reconnection occur. As the string moves forward, multiple such points of magnetic reconnection can happen. The shocks produced by these multiple points of reconnection will move outward. These shocks will have different velocities and the collision of these relativistic shocks will result in the generation of a GRB.
We consider a model where a slow shock is followed by a faster shock so that we can calculate the total energy of the plasma at the time the two shocks collide and merge with each other. Though there may be multiple shocks, for simplicity, we consider the case of only two shocks colliding with one another. We calculate an approximate value of the energy that is generated due to the collision of the shocks and the burst duration. Our basic assumption is that the shocks have different relativistic velocities and hence different $\gamma$ values. We have assumed that the first shock is the slower shock and has a relativistic velocity given by $\gamma_{1}v_{1}$ and the second shock is the faster shock having a relativistic velocity $\gamma_{2}v_{2}$. After the collision, the two shocks merge together and move as one with the merge density $\rho_{m}$ and merge Lorentz factor $\gamma_{m}$. The total kinetic energy released by the merger of the shocks is given by,
$K.E=\frac{1}{2}\rho_{m}\gamma_{m}^{2}c^{2}\frac{B}{(4\pi\rho)^{1/2}\sigma l}.$ (1)
For different models of reconnection, the length of the diffusion region $l$ is what makes a difference in the rate of reconnection and the energy emitted Kulsrud ; yamada_review . In the SP model, it value is over the whole region where reconnection occurs whereas in the other models it is scaled due to the emergent shocks which increases the overall reconnection rate. The time duration of the energy released is $\Delta t=\frac{r_{col}}{2\gamma_{1}^{2}c}$ beniamani . Here, $r_{col}$ denotes the radius over which the collision takes place. This too is model dependent.
So our model will be based on the dimensions of the string wake that we are considering. In our simulations, we have considered the energy scale for the symmetry-breaking phase transition for Abelian Higgs string to be of the order of $2.2\times 10^{15}$ GeV Lizarraga_2016 . The order of $G\mu$ for such Abelian Higgs cosmic string is therefore approx $10^{-11}-10^{-8}$. There is also the current observational limit from the Planck Collaboration Planck_2016 which gives the upper limit of the value of $G\mu$ as $\sim 10^{-7}$. For the SP model, we can then estimate the length of the reconnection region for the cosmic string case. The total length of the cosmic string wake is taken to be $L$. This is the longest dimension of the wake. Reconnection however does not occur for the full length of the wake. To estimate $l$, we have to obtain the hydromagnetic velocity. This hydromagnetic velocity depends on the density of the plasma and the magnetic field in the wake region. The $l$ of the reconnection region is then obtained from,
$l=\frac{cL}{(C_{H}\sigma L)^{1/2}}.$ (2)
We have considered $c=1$ in all our calculations. Once we know the length of the reconnection region we can obtain the kinetic energy released due to the magnetic reconnection. This energy is what powers the movement of the shock waves in the plasma. So, the kinetic energy will give us the approximate velocties of the shock waves in the plasma.
III Shock-Shock collision in the wake
We initially start with two shocks moving with different velocities away from the point of magnetic reconnection. Since we are interested in the light curves emitted by the GRB, we study the energy evolution with time. We divide the whole process into three stages. In the first stage, the total energy of the plasma is the total energy of the two shock waves moving together. This is equal to the total kinetic energy of the moving shock waves. This kinetic energy however is not constant with time. As the shocks move closer to each other the total energy increases with time. For this we need to know how the velocity of the shock waves are changing with time. As the ejected velocity of the plasma following a magnetic reconnection depends on the magnetic field, we use this knowledge to determine the change in the velocity of the shocks with time. To construct a time dependent model, we write the total energy at a given time in terms of distance and time instead of the velocity. The first shock is assumed to be produced at $t=0$ and the subsequent shock is produced after a period $\Delta t$. The distance at which the collision happens is denoted by $l_{r}$. So both the shocks travel the distance $l_{r}$ from the point at which they were generated at different velocities (and therefore different times $t_{1}$ and $t_{2}$). The kinetic energy of the first shock is given by $\frac{1}{2}\rho_{1}\gamma_{1}\frac{l_{r}^{2}}{t_{1}^{2}}$ and the kinetic energy of the second shock by $\frac{1}{2}\rho_{2}\gamma_{2}\frac{l_{r}^{2}}{t_{2}^{2}}$. Before the collision, the total energy would be the sum of these two energies. At the point of collision, we obtain the maximum energy which is given by the density of the merged shocks $\rho_{m}$, the $\gamma$ factor of the merged shocks $\gamma_{m}$ and the velocity of the shock at the merger point which is again denoted by the distance and the time.
In the numerical model, the time parameter is the variable and we consider different cases for the time difference between the generation of the shock waves. Here, $\rho_{m}=(\rho_{1}+\rho_{2})/2$ and $\gamma_{m}$ is given by,
$\gamma_{m}=\frac{2~{}\sqrt{\rho_{1}^{2}\gamma_{1}^{2}+\rho_{2}^{2}\gamma_{2}^{2}}}{\rho_{1}+\rho_{2}}.$ (3)
Hence to simplify the calculations, we have considered the densities to be integral multiples (e.g, $\rho_{2}=2\rho_{1}$ etc).
Upto the point of collision, the total kinetic energy keeps increasing but after that it will start to decrease as the merged shock gradually dissipates the kinetic energy at collision. The maximum energy occurs at the point of collision. The total energy before the collision depends on time as,
$E_{1}(t)=\frac{c^{2}}{2\sqrt{4\pi\rho}\sigma l_{r}}(\rho_{1}^{2}\Gamma_{1}^{2}+\rho_{2}^{2}\Gamma_{2}^{2})t^{2}.$ (4)
At the collision of the two shocks the maximum energy is given by,
$E_{max}=\frac{\rho_{m}^{2}\Gamma_{m}^{2}c^{2}t_{col}^{2}}{2\sqrt{4\pi\rho}\sigma l_{r}}.$ (5)
Similar to ref.kobayashi we assume once the peak amplitude is obtained, the decrease in the energy is due to the angular spreading of the energy. Assuming the density will remain more or less the same, it will be the velocity which will determine the decrease in energy. As the timestep in the numerical model is small, a Taylor series expansion will result in the approximate decrease of energy with time as $\Delta E\propto(1-\frac{t_{col}^{2}}{t^{2}})$. The constant of proportionality will depend on the same parameters as the maximum energy. So after collision, the energy will decrease as $E(t)-\Delta E$. So the energy expression after the collision is given by,
$E_{2}(t)=\frac{\rho_{m}^{2}\Gamma_{m}^{2}c^{2}t_{col}^{2}}{2\sqrt{4\pi\rho}\sigma l_{r}}\left(1-(1-\frac{t_{col}^{2}}{t^{2}})\right).$ (6)
Fig.[2] shows the resulting pulse that we have obtained.
Figure 2: Change of energy density with time as two shocks waves collide and merge, followed by the decay.
We find that there is a sharp rise followed by a gentle decay. The collision is considered instantaneous and hence obtained the sharp peak. In a more realistic scenario, it would probably happen over a short period of time which will lead to a more rounded peak. For this particular pulse, we have considered the time to be at redshift $z=30$. We have calculated the background density at that temperature and time. We have also taken the average magnetic field to be the background equipartition field at that time. This has a value of $1.95\times 10^{-2}eV^{2}$. The conductivity of the plasma is about $\sigma=0.33cm^{-1}$. The length and width of the cosmic string wake determine the length of the reconnection region. At $z=30$, this turns out to be $l_{r}=1.4\times 10^{10}cm$. Initially, we do our calculations at a fixed redshift, later we check the model by varying the redshift parameter.
IV Multiple collisions and their temporal structure
The important point that we would like to make is that the magnetic reconnection in the cosmic string wake invariably results in mulitple points of magnetic reconnection. This has been illustrated in Fig[1]. So the case of the two shock collision is the simplest possible case. Multiple shocks will be generated and they may be generated from multiple points in the cosmic string wake, so there is the possibility of multiple shock collision in the cosmic string wake. However, it is important to point out that these shocks will have different velocities and therefore different $\gamma$ factors. They may not collide in the same direction. To model this kind of collisions is rather difficult. We have extended our model to include at least a few of these conditions. Extending our two shock model, we have done a simulation where multiple shocks are generated with random values of $\gamma$. These are then made to collide similar to the two shock model that we had before. We can vary the number of shocks colliding one after the other. Though not perfect we have obtained the synthetic light curve for the mulitple collisions also. This is shown in Fig.[3].
Figure 3: Energy density with time as multiple shocks collide and merge, followed by their decay. The shocks have random velocities but generally a faster shock collides with a slower shock in this figure. The number of random collisions is mentioned in the adjoining legends
In Fig.[3], we have shown the light curve for different collision rates. Here, we have assumed that there is no correlation between the multiple shocks and each shock collides only once with the shock closest to it. We have considered all the collisions between a slow shock and a faster shock. As is expected. a larger number of collisions between the shocks only increase the total energy. We also notice that it does not change the order of the total energy generated. So it appears that the number of collisions leading to the GRB will not change the GRB significantly.
We also obtain similar curves with multiple slow moving shocks. By slow moving shocks we mean that the $\gamma$ factor is less than $0.5$c. These are depicted in Fig.[4]. Here too, we get similar light curves, only now the increasing energy is less steep than in the case where one of the shocks was a fast one.
Figure 4: Energy density with time as multiple slow shocks collide and merge, followed by their decay. Here we have kept the shock velocities random but always less than $0.5$c. The number of random collisions is mentioned in the adjoining legends
In all these figures, we have not varied the density of the shocks and considered one of the shocks to have double the density of the other. As mentioned before, this was for the ease of obtaining the resultant $\gamma$ factor of the colliding shocks. As a check, we also did the shock - shock collision for different densities of the shock waves. This is then shown in Fig.[5] with the legends denoting the ratio of the densities of the shocks. As we see in this figure, we are getting different curves for the different ratios with the higher density ratios having higher energy densities.
Figure 5: Energy density with time as multiple shocks of different density ratio collide and merge, followed by their decay. The density ratio is mentioned in the adjoining legends
Finally, we have also checked the effect of varying the background magnetic field. Again this does not change the pattern but changes the overall magnitude and timescale of the GRB. The timescale of the GRB is affected as the emergence of the shocks depends on the magnetic reconnection occuring in the cosmic string wake. For our reconnection model, $\Delta t=\frac{8}{\pi}\frac{l_{r}log(S)}{C_{H}}$. Here, $S$ is the Lundquist number and $C_{H}$ depends on the magnitude of the magnetic field. This is shown in Fig.[6].
Figure 6: Energy density with time as multiple shocks collide and merge for different background magnetic fields. The corresponding background magnetic field is mentioned in the adjoining legends
A general pattern is observed for the synthetic light curve obtained from the shock-shock collision in the magnetized wakes of cosmic strings. The timescale of the collision depends upon various factors. One of the most important factor being the density of the plasma. Since we want to do the calculations for a GRB that is observable, we have chosen $z~{}=~{}30$. However GRB’s are mostly observed at much lower $z$ values. At lower $z$ values, the magnetic reconnection length, the density of the plasma as well as the magnetic field may be different. Keeping the magneitc field value to be the same, we have obtained some graphs for lower values of $z$. This is shown in Fig.[7] for multiple collisions.
Figure 7: Energy density with time as multiple shocks collide and merge for different redshift values. The corresponding values of redshift $z$ are mentioned in the adjoining legends
Since the magnetic reconnection length changes the total energy density and the collision timescales we see that we get higher intensities for lower redshift values.
V Observational possibilities of the light curves
We have theoretically modeled a single pulse of energy that can be generated from magnetic reconnection in cosmic string wakes. We have then done a superposition of such pulses for random relativistic velocities of merger to obtain a synthetic light curve for a GRB. We would like to highlight that the burst duration and the energy released from the merging shocks depends primarily on the length of the magnetic reconnection region. This in turn depends on the magnetic field lengthscale. Since the magnetic field is generated in the wake of the cosmic string, the magnetic lengthscale would depend on the length of the wake. As mentioned previously, the magnetic fields are along the $z-$direction. They are however squeezed in the $y-$direction due to the flow of the plasma around the cosmic string. The velocity perturbation in the $y-$direction comes from the conical spacetime of the cosmic string. This space time has a deficit angle which is proportional to $G\mu$. Here, the $\mu$ or mass per unit length depends on the symmetry breaking scale which generated the cosmic string in the first place. We would now like to check whether any cosmic string model would indeed generate energies in the ranges where GRB’s have been observed.
As we have mentioned before, we have considered the parameters of an Abelian Higgs cosmic string. We have calculated the reconnection length from the parameters of the wake of the cosmic string. For different values of $z$, this turns out to be different. Though our simulations are based on the energy density, one can relate the energy density to the fluence associated with a GRB. The fluence associated with these light curves can be found out by multiplying with the length associated with the wakes which are of the order of $10^{10}$ cm or less, so the fluences are found to be in the range of $\sim 10^{-5}$ erg cm-2 or less. For $z=30$, the timescale is found out to be $\Delta t=1.14$ sec. The collision time is therefore $\triangle t<2$ sec for different Lorentz factors corresponding to the two outgoing shock waves.
We have studied the available catalogue of GRBs to compare our energy scales with observed bursts for different redshifts. The table below shows the various high-redshift GRBs observed for short duration GRB’s in the BATSE satellite. This is not a exhaustive list of short duration GRB’s but a token list to show the burst duration and fluence currently being observed by various telescopes. As is seen from the table, the energy scales and the timescales both indicate that magnetic reconnection in cosmic string wakes can be considered to be the progenitor for a GRB.
No. GRBs ID Redshift Burst Duration Fluence
(z) ($T_{90}$) (sec.) (×$\times 10^{-6}$ erg cm-2)
1. 090429B 9.4 0.53 $\pm$ 0.09 0.31 $\pm$ 0.03
2. 090423 8.2 1.13 $\pm$ 0.12 0.59 $\pm$ 0.04
3. 080913 6.7 1.04 $\pm$ 0.13 0.56 $\pm$ 0.06
4. 220521A 5.6 1.81 $\pm$ 0.07 0.98 $\pm$ 0.12
5. 240419A 5.17 0.48 $\pm$ 0.22 0.08 $\pm$ 0.02
6. 090205 4.65 1.55 $\pm$ 0.31 0.19 $\pm$ 0.03
Most of our detailed calculations are done for the redshift value of $z=30$. Our motivation for this was the previously predicted signatures of cosmic string wakes in the 21 cm redshift surveys branden1 at $z=30$. Recently some studies have also been done for lower values of $z$ with the limitation of $z>15$ Oscar_2021 . For the GRB emission, the most important change due to the redshift would be the value of the background plasma density. We have found that changes in the plasma density only changes the maximum energy associalted with the light pulse. The mechanism for the generation of the GRB remains the same irrespective of the redshift. In a future work, we plan to look at actual experimental data to try and fit the model postulated in this paper at different redshifts to try and determine a signature of a cosmic string which will also support previous predicted signatures using density inhomogeneities in cosmic string wakes.
VI Summary and Conclusions
To summarize, we have worked out a simplified model for the generation of a light pulse from the collision of magnetized shocks in the wake of a cosmic string at different redshifts. Magnetic field generated in cosmic string wakes are oppositely directed on both sides of the wakes. As the string moves through the plasma, the magnetic lines are pushed together and magnetic reconnection can occur. This reconnection results in the generation of a large amount of kinetic energy. Different models of magnetic reconnection gives rise to multiple shocks. We show that the collision of these shocks can result in the generation of a strong pulse of energy. Multiple shocks colliding may give rise to a GRB light curve. Using the various paramaters of shock waves due to magnetic reconnection in cosmic string wakes we have modeled the light pulse and the synthetic light curves. We find a generic pattern for the GRB light curve. It depends primarily on the density of the background plasma and the magnetic field. There is a weak dependence on the ratio of the densities of the colliding shocks, similarly there is a weak dependency on the velocities of the colliding shocks.
There are some basic assumptions we have made for the current work. We have only considered the collision of shocks in one direction and have assumed that there is no correlation in the multiple shock mergers. This means that multiple collision will just enhance the overall energy of the outburst. A more detailed simulation would include various other factors. It is possible that after the initial collision between two shocks, the merged shock will coast for some time before colliding with another shock generated from a different point. This will require changes in the timescales and will therefore change the light curve considerably. Since we have found that the calculated fluence and the burst duration are well within the ambit of GRB’s observed by various experimental groups, we would like to improve on these simulations in a later work. Our primary objective in this work was to show conclusively that magnetic reconnection in the wakes of cosmic strings can indeed be the progenitor for a GRB.
Acknowledgment: We acknowledge funding from the DST-SERB Power Grant no.SPG/2021/002228, D.K is supported by the Power Grant no. SPG/2021.002228, of the Government of India.